# Reinforcement Learning with Elastic Time Steps

### v3

@@ Abstract
Traditional Reinforcement Learning (RL) policies are typically implemented with fixed control rates, often disregarding the impact of control rate selection. This can lead to inefficiencies as the optimal control rate varies with task requirements. We propose the Multi-Objective Soft Elastic Actor-Critic (MOSEAC), an off-policy actor-critic algorithm that uses elastic time steps to dynamically adjust the control frequency. This approach minimizes computational resources by selecting the lowest viable frequency. We show that MOSEAC converges and produces stable policies at the theoretical level, and validate our findings in a real-time 3D racing game. MOSEAC significantly outperformed other variable time step approaches in terms of energy efficiency and task effectiveness. Additionally, MOSEAC demonstrated faster and more stable training, showcasing its potential for real-world RL applications in robotics.
Index Terms:
Reinforcement Learning, AI-Based Methods, Optimization and Optimal Control
I Introduction
Model-free deep reinforcement learning (RL) algorithms have demonstrated significant value in diverse domains, from video games [1, 2] to robotic control [3, 4]. For instance, Sony’s autonomous racing cars achieved remarkable results with fixed training frequencies of 10 Hz and 60 Hz [4].
However, using suboptimal, fixed control rates presents limitations, leading to excessive caution, wasted computational resources, risky behavior, and compromised control.
Variable Time Step Reinforcement Learning (VTS-RL) has been introduced to address issues related to fixed control rates in traditional RL. By leveraging reactive programming principles [5, 6], VTS-RL performs control actions only when necessary, reducing computational load and enabling variable action durations. For example, in robotic manipulation, VTS-RL allows a robot arm to adapt its control frequency dynamically, using lower frequencies for simple tasks and higher frequencies for complex maneuvers or delicate handling [7].
Two prominent VTS-RL algorithms are Soft Elastic Actor-Critic (SEAC  [8]) and Continuous-Time Continuous-Options (CTCO) [7]. CTCO supports continuous-time decision-making with flexible option durations, enhancing exploration and robustness. However, it involves tuning multiple hyperparameters, such as radial basis functions (RBFs) and time-related parameters $\tau$ for its adaptive discount factor $\gamma$, making tuning complex in certain environments.
SEAC incorporates reward components related to task energy (number of actions) and task time, making it effective in time-constrained environments. Despite its advantages, SEAC demands careful tuning of hyperparameters to balance task, energy, and time costs to ensure optimal performance. The sensitivity of both SEAC and CTCO to hyperparameter settings presents a challenge for users aiming to fully exploit their capabilities.
@@ I Introduction
    Enhanced Training Efficiency: We show that MOSEAC achieves faster and more stable training than CTCO, highlighting its practical benefits and applicability in real-world scenarios.
The paper is organized as follows. Section II describes the current research status of variable time step RL. Section III introduces MOSEAC with its pseudocode and Lyapunov model. Section IV describes the test environments. Section V presents the simulation parameters and results. Finally, Section VI concludes the paper.
II Related Work
@@
Sharma et al. [15] proposed learning to repeat actions to mimic dynamic control rates, but this approach does not change the control frequency or reduce computational demands. Few studies have explored repetitive behaviors in real-world scenarios, such as those by Metelli et al. [16] and Lee et al. [17]. Chen et al. [18] introduced variable “control rates” using actions like “sleep,” but still involved fixed-frequency checks.
@@
Lyapunov models provide robust stability guarantees by employing a scalar function that decreases over time, ensuring system convergence to a desired equilibrium. By incorporating Lyapunov stability into our reinforcement learning framework, we ensure that the learning process remains stable, thereby preventing erratic behavior and potential system failure [19, 22].
III Multi-Objective Soft Elastic Actor and Critic
Our algorithm builds upon previous work [9], combining SEAC’s hyperparameters for balancing task, energy, and time rewards through a simple multiplication method, and applying adaptive adjustments to the remaining hyperparameters. A key improvement is the introduction of an upper limit for the hyperparameter $\alpha_{m}$. Below is an overview of the MOSEAC algorithm. This overview emphasizes the critical aspects of MOSEAC without delving into the detailed definitions. The reader can refer to [9] for details.
@@ III Multi-Objective Soft Elastic Actor and Critic
Algorithm algorithm 1 shows the pseudocode of MOSEAC. In short, MOSEAC extends the SAC algorithm by incorporating action duration $D$ into the action policy set. This expansion allows the algorithm to predict the action and its duration simultaneously. The reward is calculated using Equation 1, and its changes are continuously monitored. If the reward trend declines, $\alpha_{m}$ increases linearly at a rate of $\psi$, without exceeding $alpha_{max}$. The action and critic networks are periodically updated like the SAC algorithm based on these preprocessed rewards.
@@ 32      Perform soft-update of $\phi^{{}^{\prime}}$ and $\theta^{{}^{\prime}}$
Our algorithm employs a scalarizing strategy to optimize multiple objectives, sweeping the design space by adapting $\alpha_{m}$ during training. Unlike Hierarchical Reinforcement Learning (HRL) [24, 25] which aims for Pareto optimality [26] through layered reward policies, our approach simplifies the process. We focus on ease of use and computational efficiency, ensuring our method can easily adapt to various algorithms.
It is worth noting that the choice of $\psi$ is crucial for optimal performance, as $\psi$ represents the sweeping step of our optimization (similar to the learning parameter in gradient descent). We suggest using the pre-set $\psi$ value from our implementation to minimize the need for further adjustments. If training performance is inadequate, a high $\psi$ value might cause the reward signal’s gradient to change too quickly, leading to instability; in this case, a lower $\psi$ is recommended. Conversely, if training progresses too slowly, a low $\psi$ value might weaken the reward signal, hindering convergence; thus, increasing $\psi$ could be beneficial.
III-A Convergence Proof
@@
To analyze the impact of dynamically adjusting $\alpha_{m}$ and $\alpha_{\varepsilon}$, we assume:
  1. 1.
    Dynamic Adjustment Rules: $\alpha_{m}$ increases monotonically by a small increment $\psi$ if the reward trend decreases over consecutive episodes, and its upper limit is $\alpha_{max}$, which guarantees algorithmic convergence and prevents reward explosion. $\alpha_{\varepsilon}$ decreases as defined.
  2. 2.
    Learning Rate Conditions: $\alpha_{k}$ and $\beta_{k}$ must satisfy [27]:
@@
  1. 1.
    Positive Scaling: As $0\leq R_{\tau}\leq 1$ and $\alpha_{m}\geq 0$, $\alpha_{m}$ only scales the reward without altering its sign. This scaling does not change the direction of the policy gradient but affects its magnitude.
  2. 2.
    Small Offset: $\alpha_{\varepsilon}$ is a small constant used to accelerate training. This small offset does not affect the direction of the policy gradient but introduces a minor shift in the value function, which does not alter the overall policy update direction.
@@ III-B Lyapunov Stability
$\dot{V}(t)=\begin{cases}\begin{split}&\alpha_{m}(t)k+\sum_{s}2\left[Q(s,a,D,\alpha_{m})-Q^{*}(s,a,D)\right]\\ &\quad\dot{Q}(s,a,D,\alpha_{m}),\quad\text{if }\alpha_{m}(t)<\alpha_{\text{max}}\end{split}\\ &\\ \begin{split}&\sum_{s}2\left[Q(s,a,D,\alpha_{m})-Q^{*}(s,a,D)\right]\\ &\quad\dot{Q}(s,a,D,\alpha_{m}),\quad\text{if }\alpha_{m}(t)=\alpha_{\text{max}}\end{split}&\end{cases}$ (13)
@@
IV Experimental Setup
We validate our MOSEAC in a real-time racing game, Trackmania [11]. Figure 1 illustrates the testing environment. In Trackmania, players race to complete the track as quickly as possible. We employed the map developed by the TMRL [2] team for a direct comparison of MOSEAC’s performance against their SAC model. It is worth noting that training and deploying a policy for this game can only happen in real time [28], ensuring the realism of the overall experiments.
@@ IV Experimental Setup
The reward system in the game is designed to encourage efficient path following: more rewards are given for covering more path points in a single move.This path is a series of evenly distributed points that collectively form the shortest route from the starting point to the end point of the racing game. This approach is consistent with the TMRL team’s methodology [2]. The game uses realistic car physics, but there is no car damage or crash detection. Further details are available in Table I.
TABLE I: Trackmania Environments Details
@@
To evaluate the impact of our variable time step approach, we used the same visual-based navigation and task reward policy across the tested approaches, varying only the control rate. It is worth noting that the task reward policy only includes time and disregards collisions.
@@
V Experimental Results
We conducted experiments with MOSEAC, CTCO [7] and SEAC [8] on Trackmania for over 1320 hours [^1] Our code is publicly available on GitHub: https://github.com/alpaficia/TMRL_SEAC. These experiments were conducted on a I5-13600K computer with an NVIDIA RTX 4070 GPU. The final result video for MOSEAC is publicly available [^2] https://youtu.be/1aQ0xSK55nk, with 43.202 seconds to complete the test track.
@@ V Experimental Results
Before discussing the performance of MOSEAC and CTCO, it is essential to provide some background context. In our previous work, the SEAC model policies were trained using different hyperparameters [29]. These hyperparameters were explicitly tuned for SEAC, optimizing its performance in the given environment. Besides, the SAC model was trained by the TMRL team with a different set of hyperparameters tailored to their training approach [2]. Given these differences, directly comparing their training curves would be unfair; hence, we focus solely on the MOSEAC and CTCO training curves.
@@
Figure 4 demonstrates that MOSEAC reaches a high point at around 6 million steps, indicating initial training success. However, Figure 5 shows that after 6 million steps, MOSEAC required over 14 million additional steps to approach the optimal value. This indicates that MOSEAC and CTCO require substantial time for initial training and further optimization in a sparse reward environment.
@@
Training in the Trackmania environment is real-time, and our 1320+ hours of training have demonstrated that MOSEAC can handle complex, sparse reward environments, albeit requiring significant time investment. Therefore, if a direct comparison with our results, specifically regarding training speed within the same reward environment, is unnecessary, we recommend optimizing the reward signals in the TMRL Trackmania environment. One possible optimization strategy could be to amplify the per-step reward signal appropriately. This adjustment theoretically allows for faster training of effective control policies, significantly enhancing training efficiency and saving substantial real-world time.
Figure 6: This figure compares the energy costs of four different algorithms: MOSEAC, CTCO, SEAC, and SAC (20Hz). The y-axis represents the energy cost in steps, while the x-axis lists the algorithms. The box plots show that MOSEAC has the lowest median energy cost, followed by SEAC, CTCO, and SAC (20Hz). This indicates that MOSEAC is the most energy-efficient among the four algorithms.
We compared the racetrack time (referred to as time cost) as well as the computational energy cost (in terms of the number of control steps) of MOSEAC with SEAC, CTCO, and SAC (20Hz) after training. Figure 6 illustrates the energy cost distribution and Figure 7 the time cost (lower is better in both cases). For both energy and time there is no overlap among the different methods, and the data follows a normal distribution (Shapiro-Wilk test of normality, see Table II and Table IV).
Using a paired sample T-test, MOSEAC showed lower energy in all of the trials compared with CTCO ($t=-64.85$, $df=29$, $p\ll 0.001$), SEAC ($t=-35.90$, $df=29$, $p\ll 0.001$), and SAC ($t=-196.35$, $df=29$, $p\ll 0.001$). The statistics for the energy cost measures are presented in Table III. Similarly, MOSEAC showed lower race time in all of the trials compared with CTCO ($t=-55,67$, $df=29$, $p\ll 0.001$), SEAC ($t=-32.92$, $df=29$, $p\ll 0.001$), and SAC ($t=-41.06$, $df=29$, $p\ll 0.001$). The statistics for the time cost measures are presented in Table V.
TABLE II: Test of Normality (Shapiro-Wilk) - Energy cost difference
@@
TABLE III: Energy cost descriptives
@@
Figure 7: This figure compares the time costs of four different algorithms: MOSEAC, CTCO, SEAC, and SAC (20Hz). The y-axis represents the time cost in seconds, while the x-axis lists the algorithms. The box plots reveal that MOSEAC has the lowest median time cost, indicating faster task completion compared to the other algorithms. CTCO, SEAC, and SAC (20Hz) show higher median time costs, with MOSEAC outperforming them in terms of time efficiency.
TABLE IV: Test of Normality (Shapiro-Wilk) - Time cost difference
@@
TABLE V: Time cost descriptives
@@
The improved performance of MOSEAC over SEAC, CTCO, and SAC (20Hz) can be attributed to its reward function and the integration of a state variable. While SEAC’s reward function is linear, combining task reward, energy penalty, and time penalty independently, MOSEAC introduces a multiplicative relationship between task reward and time-related reward. This non-linear interaction enhances the reward signal, mainly when both task performance and time efficiency are high, and naturally balances these factors. By keeping the energy penalty separate, MOSEAC maintains flexibility in tuning without complicating the relationship between time and task rewards. This design allows MOSEAC to guide the agent’s decisions more effectively, resulting in better energy efficiency and task completion speed in practical applications.
VI Conclusions

### v4

@@ Abstract
Traditional Reinforcement Learning (RL) policies are typically implemented with fixed control rates, often disregarding the impact of control rate selection. This can lead to inefficiencies as the optimal control rate varies with task requirements. We present the Multi-Objective Soft Elastic Actor-Critic (MOSEAC), an off-policy actor-critic algorithm that uses elastic time steps to dynamically adjust the control frequency. This technique minimizes computational resources by selecting the lowest viable frequency. We demonstrate that MOSEAC converges and produces stable policies at the theoretical level, and validate our findings in a real-time 3D racing game. MOSEAC significantly outperformed other variable time step approaches in terms of energy efficiency and task effectiveness. Additionally, MOSEAC shown faster and more stable training, showcasing its potential for real-world RL applications in robotics.
keywords:
Reinforcement Learning , Learning and Adaptive Systems , Optimization and Optimal Control
1 Introduction
Model-free deep Reinforcement Learning (DRL) has shown notable value in diverse domains, from video games [23, 25] to robotic control [10, 29]. For example, Sony’s autonomous racing cars achieved remarkable results with fixed training frequencies of 10 Hz and 60 Hz [29].
However, suboptimal fixed control rates present limitations, leading to excessive caution, wasted computational resources, risky behavior, and compromised control.
Variable Time Step Reinforcement Learning (VTS-RL) has been introduced to address these issues in traditional RL by utilizing reactive programming principles [16, 4]. VTS-RL performs control actions only when necessary, reducing computational load and enabling variable action durations. For example, in robotic manipulation, VTS-RL allows a robot arm to dynamically adjust its control frequency, utilizing lower frequencies for simple tasks and higher frequencies for complex maneuvers or delicate handling [11].
Two notable VTS-RL algorithms are Soft Elastic Actor-Critic (SEAC) [26] and Continuous-Time Continuous-Options (CTCO) [11]. CTCO supports continuous-time decision-making with flexible option durations, enhancing exploration and robustness. However, it requires tuning multiple hyperparameters, such as radial basis functions (RBFs) and time-related parameters $\tau$ for its adaptive discount factor $\gamma$, making tuning complex in certain environments.
SEAC incorporates reward components related to task energy (number of actions) and task time, making it effective in time-constrained environments. Despite its advantages, SEAC requires careful tuning of hyperparameters to balance task, energy, and time costs to ensure optimal performance. The sensitivity of both SEAC and CTCO to hyperparameter settings presents a challenge for users aiming to fully exploit their capabilities.
@@ 1 Introduction
    Enhanced Training Efficiency: We demonstrate that
    MOSEAC achieves faster and more stable training than CTCO, highlighting its practical benefits and applicability in real-world scenarios.
The paper is organized as follows. Section 2 describes the current research status of variable time step RL. Section 3 introduces MOSEAC with its pseudocode and Lyapunov model. Section 4 describes the test environments. Section 5 presents the simulation parameters and results. Finally, Section 6 concludes the paper.
2 Related Work
@@
Sharma et al. [21] proposed learning to repeat actions to mimic dynamic control rates, but this technique does not change the control frequency or reduce computational demands. Few studies have explored repetitive behaviors in real-world scenarios, such as those by Metelli et al. [17] and Lee et al. [13]. Chen et al. [5] introduced variable “control rates” utilizing actions like “sleep,” but still involved fixed-frequency checks.
@@
Lyapunov models provide robust stability guarantees by employing a scalar function that decreases over time, ensuring system convergence to a desired equilibrium. By incorporating Lyapunov stability into our RL system, we ensure that the learning process remains stable, thereby preventing erratic behavior and potential system failure [7, 6].
3 Multi-Objective Soft Elastic Actor and Critic
Our algorithm builds upon previous work [27], combining SEAC’s hyperparameters for balancing task, energy, and time rewards through a simple multiplication technique, and applying adaptive adjustments to the remaining hyperparameters. A key improvement is the introduction of an upper limit for the hyperparameter $\alpha_{m}$. Below is an overview of the MOSEAC algorithm. This overview emphasizes the critical aspects of MOSEAC without delving into the detailed definitions. The reader can refer to [27] for details.
@@ 3 Multi-Objective Soft Elastic Actor and Critic
Algorithm 1 presents the pseudocode for MOSEAC. MOSEAC extends the SAC algorithm by incorporating action duration $D$ into the action policy set, enabling it to predict both the action and its duration simultaneously. The reward is computed utilizing Equation 1 and is continuously monitored. If the reward trend declines, $\alpha_{m}$ increases linearly at a rate of $\psi$, but does not exceed $\alpha_{max}$. The action and critic networks are periodically updated, similar to the SAC algorithm, based on these preprocessed rewards.
@@ 32      Perform soft-update of $\phi^{{}^{\prime}}$ and $\theta^{{}^{\prime}}$
Our algorithm employs a scalarizing strategy to optimize multiple objectives, sweeping the design space by adapting $\alpha_{m}$ during training. Unlike Hierarchical Reinforcement Learning (HRL) [8, 14] which aims for Pareto optimality [18] through layered reward policies, our technique simplifies the process. We focus on ease of use and computational efficiency, ensuring our technique can easily adapt to various algorithms.
It is important to note that the choice of $\psi$ is crucial for optimal performance, as $\psi$ represents the sweeping step of our optimization (similar to the learning parameter in gradient descent). We suggest utilizing the pre-set $\psi$ value from our implementation to minimize the need for further adjustments. If training performance is inadequate, a high $\psi$ value might cause the reward signal’s gradient to change too quickly, leading to instability; in this case, a lower $\psi$ is recommended. Conversely, if training progresses too slowly, a low $\psi$ value might weaken the reward signal, hindering convergence; thus, increasing $\psi$ could be beneficial.
3.1 Convergence Proof
@@
To analyze the impact of dynamically adjusting $\alpha_{m}$ and $\alpha_{\varepsilon}$, we assume the following conditions. The dynamic adjustment rules specify that $\alpha_{m}$ increases monotonically by a small increment $\psi$ if the reward trend decreases over consecutive episodes. Its upper limit, $\alpha_{max}$, guarantees algorithmic convergence and prevents reward explosion. On the other hand, $\alpha_{\varepsilon}$ decreases as defined.
The learning rate conditions must be satisfied, where $\alpha_{k}$ and $\beta_{k}$ adhere to the following equations [12]:
@@
The positive scaling condition states that as $0\leq R_{\tau}\leq 1$ and $\alpha_{m}\geq 0$, $\alpha_{m}$ only scales the reward without altering its sign. This scaling does not change the direction of the policy gradient but affects its magnitude. Additionally, a small offset $\alpha_{\varepsilon}$ is used to accelerate training. This small offset does not affect the direction of the policy gradient but introduces a minor shift in the value function, which does not alter the overall policy update direction.
@@ 3.2 Lyapunov Stability
$\dot{V}(t)=\begin{cases}\begin{aligned} &\alpha_{m}(t)k+\sum_{s}2\left[Q(s,a,D,\alpha_{m})-Q^{*}(s,a,D)\right]\\ &\quad\dot{Q}(s,a,D,\alpha_{m}),\quad\text{if }\alpha_{m}(t)<\alpha_{\text{max}}\end{aligned}\\ &\\ \begin{aligned} &\sum_{s}2\left[Q(s,a,D,\alpha_{m})-Q^{*}(s,a,D)\right]\\ &\quad\dot{Q}(s,a,D,\alpha_{m}),\quad\text{if }\alpha_{m}(t)=\alpha_{\text{max}}\end{aligned}&\end{cases}$ (13)
@@
4 Experimental Setup
We validate our MOSEAC in a real-time racing game, Trackmania [1]. Figure 1 illustrates the testing environment. In Trackmania, players race to complete the track as quickly as possible. We employed the map developed by the TMRL [25] team for a direct comparison of MOSEAC’s performance against their SAC model. It is important to note that training and deploying a policy for this game can only happen in real time [20], ensuring the realism of the overall experiments.
@@ 4 Experimental Setup
The reward system in the game is designed to encourage efficient path following: more rewards are given for covering more path points in a single move.This path is a series of evenly distributed points that collectively form the shortest route from the starting point to the end point of the racing game. This technique is consistent with the TMRL team’s methodology [25]. The game uses realistic car physics, but there is no car damage or crash detection. Further details are available in Table 1.
Table 1: Trackmania Environments Details
@@
To evaluate the impact of our variable time step technique, we used the same visual-based navigation and task reward policy across the tested approaches, varying only the control rate. It is important to note that the task reward policy only includes time and disregards collisions.
@@
5 Experimental Results
We conducted experiments with MOSEAC, CTCO [11] and SEAC [26] on Trackmania for over 1320 hours [^2] Our code is publicly available on GitHub: https://github.com/alpaficia/TMRL_MOSEAC. These experiments were conducted on a I5-13600K computer with an NVIDIA RTX 4070 GPU. The final result video for MOSEAC is publicly available [^3] https://youtu.be/1aQ0xSK55nk, with 43.202 seconds to complete the test track.
@@ 5 Experimental Results
Before discussing the performance of MOSEAC and CTCO, it is essential to provide some background context. In our previous work, the SEAC model policies were trained utilizing different hyperparameters [28]. These hyperparameters were explicitly tuned for SEAC, optimizing its performance in the given environment. Besides, the SAC model was trained by the TMRL team with a different set of hyperparameters tailored to their training technique [25]. Given these differences, directly comparing their training curves would be unfair; hence, we focus solely on the MOSEAC and CTCO training curves.
@@
Figure 4 demonstrates that MOSEAC reaches a high point at around 6 million steps, indicating initial training success. However, Figure 5 shows that after 6 million steps, MOSEAC required over 14 million additional steps to technique the optimal value. This indicates that MOSEAC and CTCO require substantial time for initial training and further optimization in a sparse reward environment.
@@
Training in the Trackmania environment is real-time, and our 1320+ hours of training have shown that MOSEAC can handle complex, sparse reward environments, albeit requiring notable time investment. Therefore, if a direct comparison with our results, specifically regarding training speed within the same reward environment, is unnecessary, we recommend optimizing the reward signals in the TMRL Trackmania environment. One possible optimization strategy could be to amplify the per-step reward signal appropriately. This adjustment theoretically allows for faster training of effective control policies, significantly enhancing training efficiency and saving substantial real-world time.
Figure 6: This figure compares the energy costs of four different algorithms: MOSEAC, CTCO, SEAC, and SAC (20Hz). The y-axis represents the energy cost in steps, while the x-axis lists the algorithms. The box plots demonstrate that MOSEAC has the lowest median energy cost, followed by SEAC, CTCO, and SAC (20Hz). This indicates that MOSEAC is the most energy-efficient among the four algorithms.
We compared the racetrack time (referred to as time cost) as well as the computational energy cost (in terms of the number of control steps) of MOSEAC with SEAC, CTCO, and SAC (20Hz) after training. Figure 6 illustrates the energy cost distribution and Figure 7 the time cost (lower is better in both cases). For both energy and time there is no overlap among the different methods, and the data follows a normal distribution (Shapiro-Wilk test of normality, see Table 2 and Table 4).
Using a paired sample T-test, MOSEAC showed lower energy in all of the trials compared with CTCO ($t=-64.85$, $df=29$, $p\ll 0.001$), SEAC ($t=-35.90$, $df=29$, $p\ll 0.001$), and SAC ($t=-196.35$, $df=29$, $p\ll 0.001$). The statistics for the energy cost measures are presented in Table 3. Similarly, MOSEAC showed lower race time in all of the trials compared with CTCO ($t=-55,67$, $df=29$, $p\ll 0.001$), SEAC ($t=-32.92$, $df=29$, $p\ll 0.001$), and SAC ($t=-41.06$, $df=29$, $p\ll 0.001$). The statistics for the time cost measures are presented in Table 5.
Table 2: Test of Normality (Shapiro-Wilk) - Energy cost difference
@@
Table 3: Energy cost descriptives
@@
Figure 7: This figure compares the time costs of four different algorithms: MOSEAC, CTCO, SEAC, and SAC (20Hz). The y-axis represents the time cost in seconds, while the x-axis lists the algorithms. The box plots reveal that MOSEAC has the lowest median time cost, indicating faster task completion compared to the other algorithms. CTCO, SEAC, and SAC (20Hz) demonstrate higher median time costs, with MOSEAC outperforming them in terms of time efficiency.
Table 4: Test of Normality (Shapiro-Wilk) - Time cost difference
@@
Table 5: Time cost descriptives
@@
The improved performance of MOSEAC over SEAC, CTCO, and SAC (20Hz) can be attributed to its reward function and the integration of a state variable. SEAC uses a linear reward function that independently combines task reward, energy penalty, and time penalty. In contrast, MOSEAC employs a multiplicative relationship between task and time-related rewards. This non-linear interaction enhances the reward signal, especially when task performance and time efficiency are high, thereby naturally balancing these factors. By keeping the energy penalty separate, MOSEAC retains flexibility in tuning without complicating the relationship between time and task rewards. This design allows MOSEAC to guide the agent’s decisions more effectively, resulting in improved energy efficiency and faster task completion in practical applications.
6 Conclusions